OpenStudy (anonymous):

simplify sqr(500x^5 y^7) please help

OpenStudy (ash2326):

We have \[\sqrt{500x^5 y^7}\] @AbbeyD16 tell me what's \(\sqrt {a^2}\)?

OpenStudy (anonymous):

dont know :s

OpenStudy (ash2326):

do you know how to find root?

OpenStudy (anonymous):

no

OpenStudy (ash2326):

just a minute

OpenStudy (anonymous):

ok

OpenStudy (ash2326):

okay, do you know the square of 2?? \[2^2\] ???

OpenStudy (anonymous):

yeah ok

OpenStudy (ash2326):

What is it?

OpenStudy (anonymous):

4?

OpenStudy (ash2326):

good:)

OpenStudy (ash2326):

so just you have square of a number, we also have root of a number. You squared 2 to get 4 so the root of 4 is 2, do you get my point?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

so could you tell me root of 16 \[\sqrt{16}\]??

OpenStudy (anonymous):

its 4

OpenStudy (ash2326):

good :) similarly we have square and roots of variables like square of x is x^2 so root of x^2 is?

OpenStudy (anonymous):

x?

OpenStudy (ash2326):

great:D so what's the root of x^4?

OpenStudy (anonymous):

x^2

OpenStudy (ash2326):

Till now, we had easy cases. But sometimes roots are not found directly like 8 \[8=2^3\] here root is not easy \[\sqrt 8=\sqrt{2^2\times 2}\] do you get this?

OpenStudy (ash2326):

@AbbeyD16

OpenStudy (anonymous):

yea i got the concept :)

OpenStudy (ash2326):

great so in these cases we;ll try to simplify the powers which are multiples of 2 and leave the others in the root \[\sqrt{8}=\sqrt{2^2\times 2}=2\sqrt 2\] this is the simplified form of root 8 do you get this?

OpenStudy (anonymous):

yea i understand

OpenStudy (anonymous):

sorry my net is slow